# DESI Data Release 1: Stellar Catalogue

### v1

@@ 5 Discussion
Spectrophotometric distances for DESI stars. They are available in two independently derived VACs (SpecDis and SPDist). One is based on the neural network trained on the DESI spectra and Gaia parallaxes Li et al. (2025), and another trained on the SP stellar parameters, calibrated on the globular clusters and dwarf galaxies (Thomas & Battaglia, 2022). Both methods are able to provide distances for a large number of stars with better than 25% accuracy.
@@
The samples of BHBs, RRLs and halo stars with spectroscopic distance estimates provide a useful set of stars for chemo-dynamic studies of the MW halo, substructure search, and constraining the MW potential and Dark Matter distribution (see, e.g., Naidu et al., 2020; Dodd et al., 2023; Xue et al., 2008). Due to DESI’s depth, the BHB and RRL catalogs probe distances beyond 100 kpc (Byström et al., 2024; Medina et al., 2025b). Similarly, samples of red giant branch stars in DESI probe a large range of Galactocentric distances. If we use the distances from the VAC by Li et al. (2025) and require better than 30% distance accuracy we can estimate that MWS has more than 50,000 stars with heliocentric distances larger than 10 kpc, and more than 3,000 stars with distances larger than 50 kpc.
@@
The large number of metallicities measured by DESI for halo stars enables studies of the metallicity distribution functions of different stellar halo components and their changes throughout the Galaxy. Also, as shown in Figure 7, DESI detects a large number of metal-poor stars ( 110,000 stars with [Fe/H] and 5,000 stars with [Fe/H]). This enables not only the search for some of the most metal-poor stars in the Milky Way (Caffau et al., 2011), but also more broadly explore the properties of much larger homogeneous samples of extremely metal-poor stars than previously available (Aguado et al., 2019; Youakim et al., 2020).
@@
Aside from studying the overall stellar halo, the data from this VAC already contains a large number of stars that are members of various substructures in the Galaxy such as globular and open clusters, dwarf galaxies, and stellar streams. To illustrate that, in Table 5 we list the systems with more than 10 members or likely members that have been observed by DESI and are included in this VAC. We used the catalogues from Cantat-Gaudin et al. (2020), Baumgardt & Vasiliev (2021), Ibata et al. (2024) and Pace et al. (2022) for open clusters, globular clusters, stellar streams and dwarf galaxies, respectively. When the membership probability was available, we only considered stars with . We can see that this VAC contains more than 10 likely members for at least 16 clusters, 24 streams, and 4 dwarf galaxies888There are also several ultra-faint dwarf galaxies with potential members, such as Bootes 2, Bootes 3, Hercules 1, Leo 4, Leo 5, Pegasus 3, Ursa Major 1, and Willman 1.. Particularly notable is the large number of potential stellar stream members in DESI999We expect this number is underestimated due the incompleteness of the catalogue of Ibata et al. (2024) and new streams being discovered, such as Tian et al. (2024).. The key reason for this is the magnitude range of the bright survey and high completeness of DESI in that magnitude range. We also note that the number of open clusters with DESI data is severely underestimated, because the open cluster member catalogue Cantat-Gaudin et al. (2020) is incomplete at faint () magnitudes (Hunt & Reffert, 2023).
@@
To illustrate DESI’s efficiency in sampling streams and galaxies in the MW halo, in Figure 15 we show what would be the required absolute magnitude of a stellar system within the DESI footprint at different heliocentric distances to have more than 10 stars in the DESI catalogue. To calculate this we simulate 10 Gyr old stellar populations with different metallicities using MIST isochrones (Choi et al., 2016; Dotter, 2016) and Chabrier initial mass function (Chabrier, 2003). We place the systems at different heliocentric distances and then use the DESI DR1 spectroscopic targeting completeness of for main sequence and distant red-giant branch stars with at high galactic latitude ()101010The only case where the assumption of constant spectroscopic completeness would break down is the case of an extremely dense stellar system such as dense star cluster. We assume the systems are within the currently observed DR1 footprint above a galactic latitude of 40 degrees. Curves of different colours show magnitude limits for different metallicities of [Fe/H], and for the stellar populations. We also show the curve expected for DR2 where the spectroscopic completeness for targets within an even larger footprint on the sky is close to 40%. The change of the magnitude limit with distance is driven by changes in the number of stars in the DESI magnitude range and the feature at 50 kpc is due to the loss of the main sequence turn-off stars. We can see that with existing DESI sampling we expect to detect more than 10 stars in streams or dwarf galaxies that are brighter than (within the footprint) at 20 kpc, and brighter than at 100 kpc. This magnitude limit will shift by approximately two magnitudes for the DR2. This shows that DESI gives us a uniquely rich view of all the accreted systems in the MW halo, such as streams and faint dwarfs. We note, however, that while in our calculation we only check which systems will have more than 10 stars observed by DESI not all of these substructures may be easily separated from the background.
@@
While Table 5 gives just the number of stars in various clusters, dwarf galaxies, and stellar streams in DESI, each member star has a DESI metallicity and radial velocity measurement. The detailed analysis of these is beyond the scope of the paper, but in the Appendix D, we measure the mean abundances and intrinsic metallicity spread in each system using a mixture model and calibrated abundances from DESI. Table 7 provides the mean metallicity of each system and intrinsic metallicity spread for both SP and RVS pipelines. A more detailed analysis is coming in future papers.
@@
The DESI data have been already demonstrated to be extremely valuable for the analysis of stellar streams (Valluri et al., 2025; Aganze et al., 2025), detection of substructures in the MW halo (Kim et al., 2025), studies of the large scale kinematics of the Galactic halo (Byström et al., 2024), studies of white dwarfs (Manser et al., 2024), studies of the stellar halo in M31 (Dey et al., 2023), and detection of very metal-poor stars (Allende Prieto et al., 2023). Upcoming publications based on DR1 and DR2 data include studies of the dwarf spheroidal galaxies in DESI DR1, studies of the RRL population, and studies of the perturbations of the MW disk.
@@ 5.2 Future
At the moment of writing this article (April 2025), the DESI survey is still ongoing – under its fourth year of operation of the main survey. With the bright program of the main survey ahead of schedule and having finished the planned four passes of the survey, it is currently performing the fifth pass of the sky. The backup program is currently 70% done. The next public data release from DESI will be based on data accumulated over three years of the DESI survey (observed before Summer 2024). This release is already available to the collaboration and under internal validation. It is expected to be released in about 2 years (Spring 2027). This release will have more than 12 million stars across all the programs and will have significant processing improvements.
@@ 6 Conclusion
More than a million of stars in the catalogue have repeated measurements of radial velocities and stellar parameters over the timespan of a year, enabling variability studies.
The bright program of the DESI main survey provides highly accurate radial velocities with formal errors ranging from 0.1 to 6 km s-1 (median error of 0.6 km s-1). The systematic floor of the radial velocities was measured from repeated observations to be km s-1.
@@
This VAC contains a large number of metal-poor stars, with 60,000 stars having [Fe/H] and several thousands of stars with [Fe/H].
@@ Acknowledgements
SK acknowledges support from the Science & Technology Facilities Council (STFC) grant ST/Y001001/1. CAP acknowledges financial support from the Spanish MICIU projects PID2020-117493GB-I00 and PID2023-149982NB-I00. T.S.L. acknowledges financial support from Natural Sciences and Engineering Research Council of Canada (NSERC) through grant RGPIN-2022-04794.
@@
The value-added catalogue described in the paper is available at https://data.desi.lbl.gov/doc/releases/dr1/vac/mws/. All the plots for this paper can be reproduced using code availalbe at https://github.com/desimilkyway/dr1_vac_plots, with the data needed for reproducing the plots available at https://doi.org/10.5281/zenodo.15317972. The notebook with the tutorial on using the DR1 VAC is available at https://github.com/desimilkyway/dr1_tutorials. The radial velocity corrections for the backup program observations are available at https://doi.org/10.5281/zenodo.15469272.
@@
astropy (Astropy Collaboration et al., 2013, 2018, 2022), duckdb (Mühleisen & Raasveldt, 2025), dynesty (Speagle, 2020; Koposov et al., 2024a), ipython (Pérez & Granger, 2007), h5py (Collette et al., 2023), healpy (Zonca et al., 2019), matplotlib (Hunter, 2007), minimint (Koposov, 2023), numdifftools https://numdifftools.readthedocs.io/en/master/, numpy (Harris et al., 2020), imf https://github.com/keflavich/imf, pytorch (Paszke et al., 2019), rvspecfit (Koposov, 2019), scipy (Virtanen et al., 2020), sqlutilpy (Koposov, 2024), q3c (Koposov & Bartunov, 2006)
@@ Appendix A Main catalogue file
The file mwsall-pix-iron.fits is the main catalogue of measurements from coadded DESI spectra containing measurements from all surveys and programs. This FITS file has several extensions. They are described in detail in K24, but we repeat a brief summary here. The RVTAB extension contains the measurements from the RVS pipeline, such as radial velocities, stellar parameters, key target information such coordinates, primary flag, the warning flag from the RVS pipeline, the survey and program the star is coming from. The SPTAB extension contains the measurements of stellar parameters and individual elements by the SP pipeline. The FIBERMAP extension contains the targeting information, such as coordinates, targeting bitmasks such as MWS_TARGET, Legacy survey photometry, some Gaia information used for targeting (partially based on Gaia DR2 and partially on DR3). The extension also contains information about the number of coadded exposures and total exposure time. The GAIA extension contains the crossmatch with Gaia DR3 with all the columns from the gaia_source table. The SCORES extension contain various quality assurance quantities on spectra the exposures as wel as information necessary to compute effective exposure times. Each extension has the same number of rows (6,372,607). Each extension is row matched to another, such that i-th row in every extension refers to the same spectrum.
@@ Appendix B Known issues
Here we list the known issues that are relevant for the users of this VAC and generally stellar spectra.
@@
The radial velocities of stars in the backup program are affected by wavelength calibration systematic biases (see Appendix C). We provide the lookup table to approximately correct these biases.
@@
The distribution of stellar parameters measured by the RVS pipeline shows a deficit of stars, likely caused by bias for stars are the top of the red giant branch.
@@
Appendix C Radial velocity systematics for the backup program in DR1
@@
Appendix D Metallicity distributions of substructures in DESI
@@
with the following priors for the background fraction, for the [Fe/H] scatter of the background stars, for the [Fe/H] scatter of the substructure, for the mean [Fe/H] of the background and for the mean [Fe/H] of the substructure. The posteriors were sampled using dynesty nested sampling code (Speagle, 2020; Koposov et al., 2024a). When we evaluate the likelihood we take into account the individual [Fe/H] uncertainties reported by the pipelines. Table 7 gives the measurement for mean metallicities and intrinsic scatter of [Fe/H] for both SP an RVS pipelines. In a handful of cases there were fewer than 10 stars satisfying the quality cuts; in such cases we do not provide the measurement. We also inspected the fitted [Fe/H] distributions and in the cases of LMS-1 and Leiptr, the metallicity distributions were very broad and lacked a clear peak, so we do not provide the measurement for these systems as a precaution.
The mean metallicities for the globular clusters observed by DESI are in good (0.1–0.2 dex) agreement with the mean metallicities reported in the literature by Harris (1996). The mean metallicities for the dwarf galaxies, on the other hand, are systematically 0.2–0.3 dex lower than those reported by Simon (2019). We believe (see Ding, 2025, in prep) that the reason for this is that DESI’s samples are typically more extended than existing spectroscopic samples and thus are probing metal-poorer outer parts of dwarfs. As expected, the table shows the abundance spreads that are typically lower for GCs and higher for dwarfs (Willman & Strader, 2012). There are, however, few cases where the measured [Fe/H] scatter for a GC is unrealistically high (0.2 dex, i.e. M 3). Also for a few systems (e.g. M 92 and Sextans) the RVS abundances show higher measured intrinsic scatter compared to SP. These discrepancies likely indicate either additional systematics in the calibrated [Fe/H] abundances or an imperfect mixture model. We leave the investigation of that for future papers. In general, we find good agreement between mean [Fe/H] measured by the RVS and SP pipelines, with the worst disagreement for NGC 5634 of 0.21 dex.

### v3

@@ 5.1 Science with the DR1 stellar sample
Spectrophotometric distances for DESI stars. They are available in two independently derived VACs (SpecDis and SPDist). One is based on the neural network trained on the DESI spectra and Gaia parallaxes Li et al. (2025), and another trained on the SP stellar parameters, calibrated on the globular clusters and dwarf galaxies (Thomas and Battaglia, 2022). Both methods are able to provide distances for a large number of stars with better than 25% precision.
@@
The samples of BHBs, RRLs and halo stars with spectroscopic distance estimates provide a useful set of stars for chemo-dynamic studies of the MW halo, substructure search, and constraining the MW potential and Dark Matter distribution (see, e.g., Naidu et al., 2020; Dodd et al., 2023; Xue et al., 2008). Due to DESI’s depth, the BHB and RRL catalogs probe distances beyond 100 kpc (Byström et al., 2025; Medina et al., 2025b). Similarly, samples of red giant branch stars in DESI probe a large range of Galactocentric distances. If we use the distances from the VAC by Li et al. (2025) and require better than 30% distance precision we can estimate that MWS has more than 50,000 stars with heliocentric distances larger than 10 kpc, and more than 3,000 stars with distances larger than 50 kpc.
@@
The large number of metallicities measured by DESI for halo stars enables studies of the metallicity distribution functions of different stellar halo components and their changes throughout the Galaxy. Also, as shown in Figure 7, DESI detects a large number of metal-poor stars ( 110,000 stars with [Fe/H] and 5,000 stars with [Fe/H]). This enables not only the search for some of the most metal-poor stars in the Milky Way (Caffau et al., 2011), but also exploration of the properties of much larger homogeneous samples of extremely metal-poor stars than previously available (Aguado et al., 2019; Youakim et al., 2020).
@@
Aside from studying the overall stellar halo, the data from this VAC already contains a large number of stars that are members of various substructures in the Galaxy such as globular and open clusters, dwarf galaxies, and stellar streams. To illustrate that, in Table 5 we list the systems with more than 10 members or likely members that have been observed by DESI and are included in this VAC. We used the catalogues from Cantat-Gaudin et al. (2020), Baumgardt and Vasiliev (2021), Ibata et al. (2024) and Pace et al. (2022) for open clusters, globular clusters, stellar streams and dwarf galaxies, respectively. When the membership probability was available, we only considered stars with . We can see that this VAC contains more than 10 likely members for at least 16 clusters, 24 streams, and 4 dwarf galaxies.999There are also several ultra-faint dwarf galaxies with potential members, such as Bootes 2, Bootes 3, Hercules 1, Leo 4, Leo 5, Pegasus 3, Ursa Major 1, and Willman 1. Particularly notable is the large number of potential stellar stream members in DESI.101010We expect this number is underestimated due the incompleteness of the catalogue of Ibata et al. (2024) and new streams being discovered, such as Tian et al. (2024). The key reason for this is the magnitude range of the bright survey and high completeness of DESI in that magnitude range. We also note that the number of open clusters with DESI data is severely underestimated, because the open cluster member catalogue Cantat-Gaudin et al. (2020) is incomplete at faint () magnitudes (Hunt and Reffert, 2023).
@@
To illustrate DESI’s efficiency in sampling streams and galaxies in the MW halo, in Figure 15 we show what would be the required absolute magnitude of a stellar system within the DESI footprint at different heliocentric distances to have more than 10 stars in the DESI catalogue. To calculate this we simulate 10 Gyr old stellar populations with different metallicities using MIST isochrones (Choi et al., 2016; Dotter, 2016) and Chabrier initial mass function (Chabrier, 2003). We place the systems at different heliocentric distances and then use the DESI DR1 spectroscopic targeting completeness of for main sequence and distant red-giant branch stars with at high galactic latitude ().111111The only case where the assumption of constant spectroscopic completeness would break down is the case of an extremely dense stellar system such as a dense star cluster. We assume the systems are within the currently observed DR1 footprint above a galactic latitude of 40 degrees. Curves of different colours show magnitude limits for different metallicities of [Fe/H], and for the stellar populations. We also show the curve expected for DR2 where the spectroscopic completeness for targets within an even larger footprint on the sky is close to 40%. The change of the magnitude limit with distance is driven by changes in the number of stars in the DESI magnitude range and the feature at 50 kpc is due to the loss of the main sequence turn-off stars. We can see that with existing DESI sampling we expect to detect more than 10 stars in streams or dwarf galaxies that are brighter than (within the footprint) at 20 kpc, and brighter than at 100 kpc. This magnitude limit will shift by approximately two magnitudes for the DR2. This shows that DESI gives us a uniquely rich view of all the accreted systems in the MW halo, such as streams and faint dwarfs. We note, however, that while in our calculation we only check which systems will have more than 10 stars observed by DESI, not all of these substructures may be easily separated from the background.
@@
While Table 5 gives just the number of stars in various clusters, dwarf galaxies, and stellar streams in DESI, each member star has a DESI metallicity and radial velocity measurement. The detailed analysis of these is beyond the scope of the paper, but in the Appendix E, we measure the mean abundances and intrinsic metallicity spread in each system using a mixture model and calibrated abundances from DESI. Table 7 provides the mean metallicity of each system and intrinsic metallicity spread for both SP and RVS pipelines. A more detailed analysis is coming in future papers.
@@
The DESI data have been already demonstrated to be extremely valuable for the analysis of stellar streams (Valluri et al., 2025; Aganze et al., 2025), detection of substructures in the MW halo (Kim et al., 2025), studies of the large scale kinematics of the Galactic halo (Byström et al., 2025), studies of white dwarfs (Manser et al., 2024), studies of the stellar halo in M31 (Dey et al., 2023), and detection of very metal-poor stars (Allende Prieto et al., 2023). Upcoming publications based on DR1 and DR2 data include studies of the dwarf spheroidal galaxies in DESI DR1, studies of the RRL population, and studies of the perturbations of the MW disk.
@@ 5.3 Future
As of April 2025, the DESI survey is still ongoing – in its fourth year of operation of the main survey. With the bright program of the main survey ahead of schedule and having finished the planned four passes of the survey, it is currently performing the fifth pass of the sky. The backup program is currently 70% done. The next public data release from DESI will be based on data accumulated over three years of the DESI survey (observed before Summer 2024). This release is already available to the collaboration and is under internal validation. It is expected to be released in about 2 years (Spring 2027). This release will have more than 12 million stars across all the programs and will have significant processing improvements.
@@ 6 Conclusion
More than a million stars in the catalogue have repeated measurements of radial velocities and stellar parameters from the RVS pipeline over the timespan of a year, enabling variability studies.
The bright program of the DESI main survey provides highly accurate radial velocities with formal uncertainties ranging from 0.1 to 6 km s-1 (median uncertainty of 0.6 km s-1). The systematic floor of the radial velocities was measured from repeated observations to be km s-1.
@@
This VAC contains a large number of metal-poor stars, with more than 100,000 stars having [Fe/H] and several thousands of stars with [Fe/H].
@@
The list of known limitations and data quality issues in the released catalogue and associated stellar spectra is documented in the Appendix B.
@@ Acknowledgements
SK acknowledges support from the Science & Technology Facilities Council (STFC) grant ST/Y001001/1. CAP acknowledges financial support from the Spanish MICIU projects PID2020-117493GB-I00 and PID2023-149982NB-I00. T.S.L. acknowledges financial support from Natural Sciences and Engineering Research Council of Canada (NSERC) through grant RGPIN-2022-04794. LBeS acknowledges support from CNPq (Brazil) through a research productivity fellowship, grant no. [304873/2025-0].
@@
This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement.
@@
The value-added catalogue described in the paper is available at https://data.desi.lbl.gov/doc/releases/dr1/vac/mws/. All the plots for this paper can be reproduced using code available at https://github.com/desimilkyway/dr1_vac_plots, with the data needed for reproducing the plots available at https://doi.org/10.5281/zenodo.15317972. The notebook with the tutorial on using the DR1 VAC is available at https://github.com/desimilkyway/dr1_tutorials. The radial velocity corrections for the backup program observations are available at https://doi.org/10.5281/zenodo.15469272.
@@
astropy (Astropy Collaboration et al., 2013, 2018, 2022), duckdb (Mühleisen and Raasveldt, 2025), dynesty (Speagle, 2020; Koposov et al., 2024b), ipython (Pérez and Granger, 2007), h5py (Collette et al., 2023), healpy (Zonca et al., 2019), matplotlib (Hunter, 2007), minimint (Koposov, 2023), numdifftools https://numdifftools.readthedocs.io/en/master/, numpy (Harris et al., 2020), imf https://github.com/keflavich/imf, pytorch (Paszke et al., 2019), rvspecfit (Koposov, 2019), scipy (Virtanen et al., 2020), sqlutilpy (Koposov, 2024), q3c (Koposov and Bartunov, 2006)
@@ Appendix A Main catalogue file
The file mwsall-pix-iron.fits is the main catalogue of measurements from coadded DESI spectra containing measurements from all surveys and programs. This FITS file has several extensions. They are described in detail in K24, but we repeat a brief summary here. The RVTAB extension contains the measurements from the RVS pipeline, such as radial velocities, stellar parameters, key target information such coordinates, primary flag, the warning flag from the RVS pipeline, the survey and program the star is coming from. The SPTAB extension contains the measurements of stellar parameters and individual elements by the SP pipeline. The FIBERMAP extension contains the targeting information, such as coordinates, targeting bitmasks such as MWS_TARGET, Legacy survey photometry, some Gaia information used for targeting (partially based on Gaia DR2 and partially on DR3). The extension also contains information about the number of coadded exposures and total exposure time. The GAIA extension contains the crossmatch with Gaia DR3 with all the columns from the gaia_source table. The SCORES extension contain various quality assurance quantities on spectra, exposures as well as information necessary to compute effective exposure times. Each extension has the same number of rows (6,372,607). Each extension is row matched to another, such that i-th row in every extension refers to the same spectrum.
@@ Appendix B Known issues
Here we list the known issues that are relevant for the users of this VAC and stellar spectra.
@@
The radial velocities of stars in the backup program are affected by wavelength calibration systematic biases (see Appendix D). We provide the lookup table to approximately correct these biases.
@@
The distribution of stellar parameters measured by the RVS pipeline shows a deficit of stars, likely caused by bias for stars at the top of the red giant branch.
@@
The measurements from RVS are not reliable as the majority of stars have close to zero, and only very large values are an indication of large spectral line broadening.
@@
Appendix C Validation of and individual element abundances
Similar to the discussion of [Fe/H] validation in Section 4.2, here we provide the comparison of measured by both SP and RVS pipelines and individual element abundances of Ca and Mg measured by SP with high-resolution literature values. Figures 17 and 18 are analogous to Figures 8 and 9 but for measurements from both RVS and SP pipelines compared to APOGEE and Gaia values. We use the same set of selection cuts as for the [Fe/H] validation described in Section 4.2.
This comparison shows that the RVS values have a systematic offset of dex, with a smaller offset of dex for the SP pipeline measurements. The scatter of the residuals is dex for both pipelines when comparing to APOGEE data. For the Gaia gspspec data, SP shows much better agreement with a scatter of 0.06 dex. The distribution of residuals between DESI and APOGEE measurements shows signs of bimodality and extended tails, indicative that DESI systematic errors are likely correlated with other stellar properties like [Fe/H], or . We also remark that the validation sample is quite metal-rich with median for APOGEE and Gaia, hence it is difficult to assess how representative the validation of is for more metal-poor stars.
Figures 19 and 20 show the comparison of [Mg/H] and [Ca/H] measured by the SP pipeline with APOGEE and GALAH measurements. These abundances are stored in the ELEM column of the SP measurement table. As the ELEM column stores the abundances of multiple elements as an array, Ca is the 2nd and Mg is 4th elements in the array. The comparison shows that Calcium measurements (green curves in Figure 20) are the most robust, with median offset with respect to GALAH and APOGEE of dex and typical scatter with respect to high-resolution measurement of 0.1 dex. The Mg measurements show signs of bimodality of the residuals (see red curves in Figure 20 and right columns in Figure 19) and also show significant systematic offset of dex with respect to APOGEE measurements and scatter of up to dex.
Appendix D Radial velocity systematics for the backup program in DR1
@@
Appendix E Metallicity distributions of substructures in DESI
@@
with the following priors for the background fraction, for the [Fe/H] scatter of the background stars, for the [Fe/H] scatter of the substructure, for the mean [Fe/H] of the background and for the mean [Fe/H] of the substructure. The posteriors were sampled using dynesty nested sampling code (Speagle, 2020; Koposov et al., 2024b). When we evaluate the likelihood we take into account the individual [Fe/H] uncertainties reported by the pipelines. Table 7 gives the measurement for mean metallicities and intrinsic scatter of [Fe/H] for both SP and RVS pipelines. In a handful of cases there were fewer than 10 stars satisfying the quality cuts; in such cases we do not provide the measurement. We also inspected the fitted [Fe/H] distributions and in the cases of LMS-1 and Leiptr, the metallicity distributions were very broad and lacked a clear peak, so we do not provide the measurement for these systems as a precaution.
The mean metallicities for the globular clusters observed by DESI are in good (0.1–0.2 dex) agreement with the mean metallicities reported in the literature by Harris (1996). The mean metallicities for the dwarf galaxies, on the other hand, are systematically 0.2–0.3 dex lower than those reported by Simon (2019). We believe (see Ding et al., 2025) that the reason for this is that DESI’s samples are typically more extended than existing spectroscopic samples and thus are probing metal-poorer outer parts of dwarfs. As expected, the table shows the abundance spreads that are typically lower for GCs and higher for dwarfs (Willman and Strader, 2012). There are, however, few cases where the measured [Fe/H] scatter for a GC is unrealistically high (0.2 dex, i.e. M 3). Also for a few systems (e.g. M 92 and Sextans) the RVS abundances show higher measured intrinsic scatter compared to SP. These discrepancies likely indicate either additional systematics in the calibrated [Fe/H] abundances or an imperfect mixture model. We leave the investigation of that for future papers. In general, we find good agreement between mean [Fe/H] measured by the RVS and SP pipelines, with the worst disagreement for NGC 5634 of 0.21 dex.